OpenStudy (anonymous):

NEED HELP AM NEW TO THIS. show that each of the following statements is an identity by transforming the left side of each one into the right side. cos θ tan θ = sin θ

Parth (parthkohli):

Ah! Are you familiar with the term known as 'isolation'? You must isolate \(\cos\theta, \sin\theta , \tan \theta\) to get different identities.

OpenStudy (anonymous):

You need to show that they are the same thing. \[\cos \theta \tan \theta = \sin \theta\] \[\cos \theta * \sin \theta/\cos \theta = \sin \theta\] the cos theta cancels and then you're left with: \[\sin \theta = \sin \theta\]

OpenStudy (anonymous):

My recommendation for problems like this is to always, ALWAYS put everything in terms of sine and cosine.